OpenStudy (kayc):

solve the inequality: (5x + 2)(x^2-1) / x + 2 > 0

OpenStudy (raden):

what zeros for all terms ?

OpenStudy (kayc):

@RadEn Sorry, i don't understand what you're asking D:

OpenStudy (raden):

ok, nopes.. first, because x^2-1=(x+1)(x-1), so it can be (5x + 2)(x+1)(x-1)/(x + 2) > 0 solve for x, if : 5x + 2=0, x=...? x+1=0, x=...? x-1=0, x=...? x+2=0, x=...? find of them ! then we continue next steps...